OpenStudy (anonymous):

Write the standard form of the equation of the circle with the given characteristics. Center: (0, 0); radius: 4

myininaya (myininaya):

(x-h)^2+(y-k)^2=r^2 (h,k)=(0,0) radius=r=4

myininaya (myininaya):

just plug in